find the radius of a circle which is concentric with the circle xsq+ysq-2x+6y-3=0 and whose radius is double the radius of the circle xsq+ysq-6x-1=0

Grade:11

1 Answers

DilipReddy Ask IITians expert-IITB
13 years ago

Dear Ramesh

As you are asking for the radius of a circle whose radius is double the radius of the circle xsq+ysq-6x-1=0.We shall first find the radius of the above circle.

It can be written as (x-3)^2+y^2-9-1=0

(x-3)^2+y^2=10

comparing the above equatio with (x-a)^2+(y-b)^2=r^2.

we get radius =√(10).

and the desired circle radius =2√(10)

centre of xsq+ysq-2x+6y-3=0 is

(x-1)^2+(y-(-3))^2=13

centre is (1,-3)

Seems you are asking for the circle equation with centre (1,-3) and radius 2√(10) is (x-1)^2+(y+3)^2=40.

 

 

thanks